OpenStudy (anonymous):

Solve x2 + 15x = -56. (5 points) (-3 (+/-) i sqrt41 )/ 4 (-3 (+/-) i sqrt55 )/ 4 (-3 (+/-) i sqrt21 )/ 4 (-3 (+/-) i sqrt15 )/ 4

OpenStudy (anonymous):

\(\cal\Huge Welcome~To~Openstudy!\)

OpenStudy (anonymous):

Thanks! Can you help me with this question?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

I don't really know where to begin solving this

OpenStudy (anonymous):

@iGreen @jojo4eva @TheSmartOne @confluxepic @Cosmichaotic

OpenStudy (anonymous):

@BrooklynRoyals I can't figure it out, I don't know where to start either.

OpenStudy (anonymous):

\(x^2-15x=-56\) <=> \(x^2-15x+56=0\) <=> \((x-7)(x-8)=0\) So x = 7 or x = 8. But I don't see the answer in the options.